OpenStudy (anonymous):

The sales tax in the city of Taxemalot is 8.1%. What is the growth factor for any item purchased at your local Tarmart? a.8.1 c.1.081 b.0.81 d.81

jimthompson5910 (jim_thompson5910):

The growth factor is simply the multiplier you multiply the initial cost by for example, say the item is $1000 you would multiply the growth factor by 1000 to get the total amount you pay (with taxes included)

jimthompson5910 (jim_thompson5910):

you find the growth rate by using the formula 1 + r r = tax rate (as a decimal)

OpenStudy (anonymous):

Ahh i see, thank you. The lecture thing that instructed me did not mention anything of the sort

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

it would be B right, i had originally thought at. As it is a decimal of its amount?

jimthompson5910 (jim_thompson5910):

1 + r = 1 + 0.081 = 1.081, correct

jimthompson5910 (jim_thompson5910):

so if the item initially costs $1000, then 1000*1.081 = 1081 is the final cost (with taxes)

OpenStudy (anonymous):

Ty again that makes much more sense. This is actually the easier side of my class but none of the online lectures explain how things work, left to my own accords lol

jimthompson5910 (jim_thompson5910):

sure thing